# SPEAR: A Unified SSL Framework for Learning Speech and Audio Representations

### v1

@@ Abstract
Self-Supervised Learning (SSL) excels at learning generic representations of acoustic signals, yet prevailing methods remain domain-specific, tailored to either speech or general audio, hindering the development of a unified representation model with a comprehensive capability over both domains. To address this, we present SPEAR (SPEech and Audio Representations), the first SSL framework to successfully learn unified speech and audio representations from a mixture of speech and audio data. SPEAR proposes a unified pre-training objective based on masked prediction of fine-grained discrete tokens for both speech and general audio. These tokens are derived from continuous speech and audio representations using a Multi-codebook Vector Quantisation (MVQ) method, retaining rich acoustic detail essential for modelling both speech and complex audio events. SPEAR is applied to pre-train both single-domain and unified speech-and-audio SSL models. Our speech-domain model establishes a new state-of-the-art on the SUPERB benchmark, a speech processing benchmark for SSL models, matching or surpassing the highly competitive WavLM Large on 12 out of 15 tasks with the same pre-training corpora and a similar model size. Crucially, our unified model learns complementary features and demonstrates comprehensive capabilities across two major benchmarks, SUPERB and HEAR, for evaluating audio representations. By further scaling up the model size and pre-training data, we present a unified model with 600M parameters that excels in both domains, establishing it as one of the most powerful and versatile open-source SSL models for auditory understanding. The inference code and pre-trained models will be made publicly available.
@@ 1 Introduction
Auditory signals, encompassing speech and general audio, are a fundamental medium for human perception and interaction (Bregman, 1994). While speech conveys linguistic and paralinguistic information for communication, general audio provides the environmental context and sound events crucial for situational awareness. Developing a unified encoder with comprehensive capabilities across both domains is therefore a significant goal for modern AI systems (Bommasani, 2021).
Self-supervised learning (SSL) has emerged as an effective paradigm for learning generic representations in the field of speech and audio processing (Baevski et al., 2020; 2022; Huang et al., 2022; Dinkel et al., 2024). By leveraging a large amount of unlabelled data during pre-training, SSL models can achieve very high performance on tasks with limited supervised data (Chen et al., 2020b; Baevski et al., 2020). In speech processing, one dominant approach is masked token prediction (Hsu et al., 2021; Chung et al., 2021; Chiu et al., 2022). This approach involves quantisation techniques, such as k-means clustering (Hsu et al., 2021) or random projection quantisation (Chiu et al., 2022), to generate coarse-grained discrete tokens from the raw speech signal or intermediate SSL representations. Although the quantisation process discards some acoustic details, such as prosody and paralinguistic information (Xin et al., 2024), these tokens still effectively capture the phonetic content of speech (Baevski et al., 2020; Hsu et al., 2021). This makes the pre-training strategy highly suitable for the speech domain and has enabled models to achieve excellent performance on a wide range of downstream speech processing tasks (Chen et al., 2022; Yang et al., 2021).
However, such coarse-grained tokens are inadequate for general audio. In contrast to speech, general audio exhibits more irregular and complex spectro-temporal patterns (Attias & Schreiner, 1996) that are difficult to capture with coarse-grained discrete units produced by methods like k-means. While BEATs (Chen et al., 2023) adapts discrete token prediction for audio, it requires a complex iterative process to train a dedicated audio tokeniser, substantially increasing training complexity. Consequently, most SSL models for general audio forgo token prediction in favour of other objectives, such as masked autoencoder (MAE) (Huang et al., 2022; Dinkel et al., 2024) or bootstrapping-based methods (Li et al., 2024a; Chen et al., 2024). This divergence in pre-training objectives has prevented the emergence of a unified SSL approach for both the speech and general audio domains.
To address this lack of unification, we propose SPEAR (SPEech and Audio Representations), a unified SSL framework based on masked prediction of fine-grained discrete tokens for both speech and general audio. Our approach applies multi-codebook vector quantisation (MVQ) (Guo et al., 2023) to the intermediate representations of existing speech and audio SSL models to obtain fine-grained discrete tokens, which then serve as the targets for a masked prediction objective. MVQ decomposes the representation space into multiple subspaces spanned by parallel codebooks. This multi-codebook design enables MVQ tokens to retain far more detail than coarse-grained discrete units, making SPEAR suitable for both speech and general audio. The pre-training objective can be viewed as performing multiple masked language modelling (Devlin et al., 2019) tasks simultaneously, one for each MVQ codebook. A key aspect of SPEAR is the joint pre-training on speech and audio, where the model learns to predict two sets of MVQ tokens derived from separate expert models from two domains. Together with a specially designed asymmetrical pre-training strategy, the dual-target objective enables SPEAR to learn a single, unified representation space bridging both domains. Finally, to enhance the model’s versatility for tasks requiring different temporal granularities (Shi et al., 2024), we integrate a multi-temporal resolution encoder (Yao et al., 2024), allowing the model to process the input signal at variable frame rates in intermediate layers.
@@
We propose SPEAR, the first unified SSL framework for both speech and general audio that successfully learns high-quality unified representations for both domains.
We conduct extensive experiments and validate the effectiveness of SPEAR in both single-domain and unified settings. Notably, our speech-domain model achieves the same or better performance than the competitive WavLM Large (Chen et al., 2022) on 12 out of 15 SUPERB (Yang et al., 2021; Tsai et al., 2022) tasks under a fair comparison. Our audio model approaches the best-performing SSL models on the HEAR (Turian et al., 2022) benchmark, and even outperforms them on environment-related tasks while using far less data.
We demonstrate that SPEAR unifies joint speech and audio pre-training, resulting in a model with a comprehensive capability over both domains. Furthermore, we show that this synergy is enhanced when scaling up model parameters and pre-training data.
Inference code and pre-trained models will be made open-source to facilitate future research.
@@ 2 Related Work
SSL for Speech and Audio
As introduced in Section 1, SSL has been widely adopted for learning generic representations in both the speech domain (Baevski et al., 2020; Hsu et al., 2021; Chen et al., 2022) and audio domain (Huang et al., 2022; Dinkel et al., 2024). Yet, existing SSL methods are typically domain‑specific, and a unified framework for both domains remains absent. Bootstrapping approaches (Grill et al., 2020; Caron et al., 2021; Niizumi et al., 2021) have shown promise, achieving strong performance in speech (Baevski et al., 2022; 2023) and audio (Chen et al., 2024; Li et al., 2024a) independently, but have not been successfully applied to joint SSL on both domains. While Gong et al. (2022) explored using both speech and audio data for SSL, their main focus was to improve general audio capability. To the best of our knowledge, our proposed SPEAR framework is the first to establish a unified SSL pre-training framework for both speech and general audio.
Masked-Token-Prediction-based SSL
Masked-token prediction is a widely used pretext task in SSL (Devlin et al., 2019). In the field of speech and audio, prevailing methods rely on coarse-grained acoustic tokens generated via k-means clustering (Hsu et al., 2021; Chung et al., 2021) or random-projection quantisation (Chiu et al., 2022; Chen et al., 2023). In the music domain, MERT (Li et al., 2024b) integrates fine-grained Encodec tokens (Défossez et al., 2023) as its pre-training targets. Our SPEAR framework likewise utilises fine-grained tokens as the pre-training target, but derives them via a non-hierarchical multi-codebook vector quantisation method.
Knowledge Distillation
Knowledge distillation (KD) (Hinton et al., 2014) is frequently employed as a model compression technique (Jiao et al., 2020; Chang et al., 2022; Yang et al., 2022). SPEAR is related to multi-teacher KD, since its pre-training targets are obtained from two domain-specific teacher models. Multi-teacher KD has been explored to combine knowledge from multiple domains within a single model (Ranzinger et al., 2024). In the speech and audio domain, Yang et al. (2025); Chang et al. (2025) adopt this strategy to distil knowledge from separate, pre-trained speech and audio models. However, these methods primarily focus on feature matching, using objectives such as the L1 loss or cosine similarity to align the student representation space with that of the teacher. Our method differs from them by coupling KD with a well‑established SSL objective for representation learning, rather than explicitly matching teacher representations.
@@ 3 SPEAR
In this section, we introduce the proposed SPEAR framework. We hypothesise that a masked prediction objective can serve as a unified SSL solution for both speech and general audio, provided the discrete tokens are sufficiently fine-grained to retain critical acoustic detail from both domains. This motivates our choice of a powerful quantisation method, which is described below.
@@ 3.1 Multi-codebook Vector Quantisation
To generate fine-grained discrete targets for our masked prediction SSL objective, we employ multi-codebook vector quantisation (MVQ) (Guo et al., 2023), a trainable quantisation method originally proposed to compress high-dimensional feature vectors for storage optimisation. To the best of our knowledge, the application of MVQ in the context of SSL pre-training has never been explored. MVQ utilises parallel codebooks, each containing trainable code vectors. Given an input feature vector , MVQ encodes it into a tuple of discrete tokens, i.e. . Each token is an integer index in the range specifying which code vector to select from the -th codebook. These selected vectors can then be used to approximate the original feature vector via a direct-sum scheme (Barnes & Watkins, 1995).
@@
Intuitively, this process partitions the feature space into distinct subspaces, each governed by a corresponding codebook. The multi-codebook structure produces significantly more fine-grained representations than coarse methods like k-means, as the number of representable states grows exponentially as . Compared to other multi-codebook quantisation methods like RVQ (Défossez et al., 2023), the codebooks in MVQ are non‑hierarchical, reducing inter‑codebook correlation and making each codebook equally important. The MVQ quantiser is trained by minimising the reconstruction error, supplemented by a diversity loss to encourage uniform code usage within each codebook. For a complete description of MVQ encoding and training mechanisms, we refer readers to the original paper (Guo et al., 2023) and the extended summary in Appendix B.
3.2 Multi-Codebook Fine-grained Masked Token Prediction
3.2.1 Single Domain Pre-training
@@
The pre-training objective is to train a single student encoder , to predict fine-grained discrete tokens extracted from a pre-trained SSL teacher model in a masked-token prediction manner. Since the teacher used for generating the pre-training targets was trained without any labelled data, SPEAR is treated as an SSL approach. An illustration of the overall framework is shown in Figure 1.
@@
The student encoder consists of a frontend processor and a feature encoder (specifically, a Zipformer (Yao et al., 2024)). The frontend processor converts the raw input waveform into frame-level acoustical representations of length . A masking operation is applied to by randomly sampling a set of frames and replacing with a learnable mask embedding , creating the masked input . The feature encoder then processes to produce a sequence of contextualised representations , where .
@@
To generate the prediction targets, the same raw audio waveform is fed into the teacher model , producing a sequence of frame-level representations . We assume the teacher and student models share the same frame rate111This can be achieved by interpolating the teacher representations if the frame rates differ.. These representations are then quantised frame-by-frame using a pre-trained MVQ quantiser to produce a sequence of fine-grained discrete tokens as the pre-training targets, where .
@@
The student model is trained to predict the target tokens from the contextualised representations . The multi-codebook masked prediction loss is formulated as the sum of independent prediction losses, one for each codebook in the MVQ quantiser. Each of these losses is a cross-entropy objective calculated over all frames, with an adjustable weight for masked and unmasked frames222The effect of is investigated in Appendix G.2:
@@
where and are the loss on masked and unmasked frames for the -th codebook, respectively. is the predicted probability of the correct token at time for the -th codebook, which is computed via a softmax function over the logits from a projection matrix :
where is the projection matrix of the prediction head for the -th codebook.
3.2.2 Unified Dual-Domain Pre-training
@@
The framework is extended to dual-domain pre-training on a mixture of speech and general audio data for learning unified representations of both domains. Specifically, we employ two expert teacher models, (speech) and (general audio), along with their corresponding pre-trained MVQ quantisers, and . Note that the number of codebooks could be different for and .
For each input waveform, the teacher representations and are extracted. Two sets of fine-grained target tokens are obtained by applying the corresponding quantiser on and :
@@
During dual-domain pre-training, the speech tokens are used as universal prediction targets for all input data, whereas the audio tokens are only used for loss computation when the input is general audio. This asymmetric approach helps the model achieve balanced performance across both domains (see Appendix G.5 for a comparison of three dual-domain pre-training strategies). The dual-domain training objective is formulated as follows:
@@
where is the single-domain masked prediction loss defined in Equation 2. The term is an indicator function that returns 1 if the input is general audio and 0 otherwise. is a hyperparameter for balancing the contribution of the general-audio-specific loss. By learning to predict both and , the student model can learn a joint feature space for both domains.

### v2

@@ Abstract
Self-supervised learning (SSL) has significantly advanced acoustic representation learning. However, most existing models are optimised for either speech or audio event understanding, resulting in a persistent gap between these two domains. We address this gap with SPEAR (SPEech and Audio Representations), a self-supervised framework that distils complementary knowledge from a speech-focused SSL teacher and a general-audio SSL teacher into a single unified model. SPEAR applies multi-codebook vector quantisation to continuous teacher representations to produce fine-grained discrete tokens that capture both semantic and acoustic information. To effectively integrate these heterogeneous representations, SPEAR jointly predicts them given a masked input with an asymmetric pre-training loss. We further improve robustness in complex sound scenes through a novel token mixing mechanism. Extensive experiments demonstrate that SPEAR consistently outperforms existing unified speech and audio models. SPEAR establishes a new state-of-the-art on the SUPERB benchmark, surpassing WavLM Large on 12 of 15 tasks, while achieving competitive performance on the HEAR benchmark. These results position SPEAR as a versatile foundation for general-purpose speech and audio representation learning. The code and pre-trained models will be released.
@@ 1 Introduction
The pursuit of general-purpose foundation models has driven transformative progress in natural language processing (Brown et al., 2020) and computer vision (Kirillov et al., 2023). In contrast, acoustic signal processing remains fragmented into three traditionally disjoint domains: speech, audio events, and music. Speech processing focuses on linguistic and paralinguistic information, typically treating non-speech sounds as noise to be suppressed. Conversely, audio event processing aims to classify acoustic events, and it often collapses speech into coarse categories, losing semantic content (e.g., “a man speaking”). Music processing lies at the intersection, requiring both acoustic event recognition for instrumentation and linguistic understanding for lyrics and prosody.
Crucially, real-world environments present complex sound scenes containing multiple, often overlapping, sound sources. Treating distinct domains in isolation neglects their intrinsic synergy: environmental sounds provide essential context for speech perception, while speech content can disambiguate sound events (Bregman, 1994). Consequently, reliance on specialised models limits holistic speech and audio understanding, motivating the need for a unified representation space. For clarity, we use the term “speech” to denote speech-centric tasks that rely on linguistic and paralinguistic information, and “audio” to collectively refer to tasks emphasising environmental sounds and music throughout the remainder of this paper.
Against this backdrop, self-supervised learning (SSL) has emerged as a powerful paradigm for learning generic representations. While its effectiveness has been demonstrated independently in speech (Baevski et al., 2020; Hsu et al., 2021; Baevski et al., 2022; Chiu et al., 2022) and audio (Chen et al., 2023; Huang et al., 2022; Pepino et al., 2025; Dinkel et al., 2024; Chen et al., 2024a) domains (Yang et al., 2021; Turian et al., 2022), constructing a single foundation model that generalises across both remains difficult due to divergent inductive biases. Speech SSL models typically utilise coarse-grained units to capture phonetic structure, which are insufficient for modelling the irregular, complex spectro-temporal patterns of general audio. Conversely, audio SSL models tend to focus on low-level acoustic details instead of high-level semantics required for speech-centric tasks. To bridge this gap, recent studies (Yang et al., 2025; Chang et al., 2025) have explored distilling complementary knowledge from domain-specific teachers into a single model.
To address this lack of unification, we propose SPEAR (SPEech and Audio Representations), a unified self-supervised learning framework for jointly learning speech and general audio representations. SPEAR is motivated by the hypothesis that masked prediction over fine-grained discrete tokens can serve as a shared learning interface for self-supervision across both domains. Concretely, SPEAR employs multi-codebook vector quantisation (MVQ) (Guo et al., 2023) to derive fine-grained discrete tokens from intermediate representations of existing speech- and audio-oriented SSL models, which are then used as supervision targets in a masked token prediction objective. Compared to conventional coarse-grained discretisation schemes (e.g., -means), MVQ preserves substantially richer acoustic and temporal detail, making the resulting tokens suitable for both speech-centric and general audio modelling. By integrating masked token prediction with knowledge distillation, SPEAR enables the student encoder to inherit complementary strengths from domain-specific teachers, while simultaneously encouraging the discovery of both high-level semantic structure and fine-grained acoustic detail directly from data. SPEAR is pre-trained on both speech and audio data, where the model jointly predicts two sets of MVQ tokens distilled from the respective domain-specific teachers. To ensure balanced learning and effective fusion across domains, we further introduce an asymmetric dual-domain objective that explicitly accounts for their differing characteristics. Finally, to improve robustness in complex sound scenes with overlapping sound sources, we propose a token mixing mechanism that constructs augmented training samples by stochastically combining supervision targets from multiple sources.
@@
We propose SPEAR, a state-of-the-art SSL framework that bridges the gap between speech and general audio representation learning by distilling complementary knowledge from domain-specific teachers into a single unified model.
We introduce a novel domain fusion strategy to integrate the heterogeneous representations from two teachers, where we jointly predict the fine-grained MVQ tokens derived from the continuous representation space of the teachers, given a masked input in an asymmetrical manner.
We propose a token mixing mechanism that stochastically fuses the targets from multiple sources, yielding consistent improvements on complex sound scenes.
@@ 2 Related Work
Speech SSL and Audio SSL
The field of speech and audio processing covers a wide range of tasks (Gold et al., 2011), such as automatic speech recognition (ASR), speaker verification (SV), music instrument classification, and audio tagging (AT). Some tasks require high-level semantic understanding (e.g., ASR), while others focus on low-level acoustic detail (e.g., AT, pitch estimation). This poses challenges in learning a unified representation space for both the speech and the general audio domains (Bengio et al., 2013). SSL has been widely adopted for learning generic representations in both the speech (Baevski et al., 2020; Hsu et al., 2021; Chen et al., 2022) and the audio domains (Huang et al., 2022; Dinkel et al., 2024). Yet, existing SSL methods are typically domain‑specific and transfer poorly to the other domain, and a unified framework for both domains remains underexplored. While bootstrapping approaches (Grill et al., 2020) have shown promise in speech (Baevski et al., 2022, 2023) and audio (Chen et al., 2024a; Li et al., 2024) independently, directly applying such objectives (e.g., data2vec) to joint pre-training on speech and audio data degrades performance by a large margin (Chang et al., 2025). Similarly, Dasheng (Dinkel et al., 2024) applied an MAE (He et al., 2022) objective to large-scale speech and audio data, but only showed limited capability on speech tasks requiring semantic understanding.
Masked-Token Prediction and Knowledge Distillation
Work most closely related to SPEAR includes EncodecMAE (Pepino et al., 2025), MT2KD (Yang et al., 2025) and USAD (Chang et al., 2025). EncodecMAE performs masked-token prediction on Encodec RVQ tokens (Défossez et al., 2023) for audio representation learning. However, it overlooks the hierarchy of RVQ tokens and is only applicable to the general audio domain, showing weak generalisation on speech tasks. In contrast, SPEAR uses non-hierarchical MVQ tokens as prediction targets and generalises to the speech domain. MT2KD (Yang et al., 2025) was the first to build a unified encoder by performing knowledge distillation (KD) on three teachers specialising in ASR, SV, and AT, using L1 and cosine similarity. However, the resulting model is constrained to the three speech and audio processing tasks since the teacher models use supervised training. USAD jointly distils two SSL teachers specialising in speech and the general audio domain into a single unified encoder. Similar to MT2KD, USAD employs feature matching losses to align the student representations with those of the teacher layer-wise. While both methods learn a unified representation space, their performance is constrained by the teacher models since they are purely mimicking the teacher representations. Compared to MT2KD and USAD, SPEAR combines KD with a well-established masked token prediction task instead of relying solely on feature matching losses. This encourages the model to learn contextualised representation by discovering semantic and acoustic structures in the data. Additionally, SPEAR considers the domain mismatch between speech and general audio, and introduces an asymmetrical pre-training loss to cope with it.
Speech and Audio Encoders for Audio LLMs
Other work related to SPEAR includes SALMONN (Tang et al., 2024), Audio Flamingo 3 (Ghosh et al., 2025) and Qwen-omni (Xu et al., 2025). SALMONN found that it is important to combine a speech specialist with an audio-event specialist for a comprehensive audio perception capability in multi-modal large language models (LLMs). Audio Flamingo 3 and Qwen-omni train an audio encoder with text LLMs end-to-end, requiring a vast amount of paired data and computation resources. However, SPEAR relies purely on unlabelled data for learning unified speech and audio representations, since the training labels are derived from domain-specific SSL teachers trained without supervision.
@@ 3.1 Multi-codebook Vector Quantisation
We hypothesise that masked-token prediction can serve as a unified SSL objective for both speech and general audio, provided the discrete tokens are sufficiently fine-grained to retain critical semantic and acoustic detail from both domains. In SPEAR, we employ multi-codebook vector quantisation (MVQ) (Guo et al., 2023), a trainable vector quantisation method to generate fine-grained discrete targets for our masked prediction SSL objective. MVQ was originally proposed to compress high-dimensional feature vectors for storage optimisation. To the best of our knowledge, the application of MVQ in the context of SSL pre-training has not been explored. MVQ utilises parallel codebooks, each containing trainable code vectors. Given an input feature vector , MVQ encodes it into a tuple of discrete tokens, i.e. . Each token is an integer index in , specifying which code to select from the -th codebook. These selected vectors can then be used to approximate the original feature vector via a direct-sum scheme (Barnes and Watkins, 1995).
@@
Intuitively, MVQ partitions a feature space into distinct subspaces, each governed by a single codebook. The multi-codebook design produces far more fine-grained representations than coarse methods, e.g., -means, as the number of representable states grows exponentially to . Compared to other multi-codebook quantisation methods such as RVQ (Défossez et al., 2023), the codebooks in MVQ are non-hierarchical and independent. MVQ is trained by minimising the reconstruction error, supplemented by a diversity loss to encourage uniform code usage within each codebook. For a complete description of MVQ encoding and training mechanisms, we refer readers to the original paper (Guo et al., 2023) and the extended summary in Appendix A.
3.2 Unified Speech and Audio Representation Learning
We first elaborate on how to leverage MVQ tokens for representation learning in a single teacher scenario, and then extend it to unified representation learning over both speech and audio domains.
3.2.1 MVQ-based Masked Token Prediction
@@
In the single teacher scenario, the pre-training objective is to train a student encoder to predict fine-grained MVQ tokens extracted from a pre-trained SSL teacher model in a masked-token prediction manner. Since the teacher used for generating the pre-training targets was trained without any labelled data, SPEAR is treated as an SSL approach. An illustration of the overall framework is shown in Figure 1.
@@
The student encoder consists of a frontend processor and a feature encoder . The frontend processor converts the raw input waveform into frame-level representations of length . A masking operation is applied to by randomly sampling a set of frames and replacing with a learnable mask embedding , creating the masked input . The feature encoder then processes to produce a sequence of contextualised representations , where .
@@
To generate the prediction targets, is also fed into the teacher model , producing a sequence of frame-level representations . We assume the teacher and student models share the same frame rate111If not, this can be achieved by interpolating the teacher representations if the frame rates differ.. These representations are then quantised frame-by-frame using a pre-trained MVQ quantiser to produce a sequence of fine-grained discrete tokens as the pre-training targets, where .
@@
The student model is trained to predict the target tokens from the contextualised representations . The multi-codebook masked prediction loss is formulated as the sum of independent prediction losses, one for each codebook in the MVQ quantiser. Each of these losses is a cross-entropy objective calculated over all frames, with an adjustable weight for masked and unmasked frames222The effect of is investigated in Appendix F.2:
@@
where and are the loss on masked and unmasked frames for the -th codebook, respectively. is the predicted probability of the correct token at time for the -th codebook, generated by independent linear prediction heads on top of the feature encoder.
3.2.2 Asymmetrical Dual-Domain Pre-training
@@
The framework is extended to dual-domain pre-training on a mixture of speech and general audio data for learning unified representations of both domains. Specifically, we employ two expert teacher models, and , along with their corresponding pre-trained MVQ quantisers, and . For each input waveform, the teacher representations and are extracted. Two sets of MVQ tokens and are obtained by applying the corresponding quantiser on the corresponding representations.
Since the teacher models are domain-specific, simply computing losses against both targets regardless of the domain of the input data could be suboptimal. Therefore, we design three dual-domain pre-training strategies:
Joint: Each sample induces two losses, computed against and , regardless of its domain.
Asymmetrical: For speech data, loss is only computed against . For audio data, loss is computed against both and .
@@
In SPEAR, the Asymmetrical strategy was adopted because it yields the best performance (see Section 5.4 for a comparison of the three strategies). That said, the speech tokens are used as universal prediction targets for all input data, whereas the audio tokens are only used as targets for general audio input. The asymmetrical dual-domain pre-train objective is formulated as follows:
@@
where is the single-domain masked prediction loss defined in Eqn. (1). The term is an indicator function that returns 1 if the input is general audio and 0 otherwise. is a hyperparameter for balancing the contribution of the general-audio-specific loss.
3.2.3 Token mixing
Prior work (Chen et al., 2022, 2024b) performs a speech denoising task by only predicting the primary target while filtering out the secondary signal in a mixed training sample, which has been shown to improve the quality of speech representations. However, we argue that treating the secondary signals merely as “noise” to suppress may degrade the model’s generality on complex sound scenes such as overlapped or noisy speech.
To address this, we introduce token mixing, a mechanism that dynamically constructs augmented training targets by stochastically combining the MVQ tokens from multiple sources based on the individual source energy for mixed audio samples. Let be the original signal and be a randomly sampled signal, we mix them at their full length with a random delay to create an augmented training sample. The clean teacher targets and of both signals are also mixed based on their signal power, forming an augmented pre-training target :
where is the mixing delay (in frames) and is a scalar mixing coefficient derived from the signal power () of the two mixed signals:
An illustration of the token mixing process is shown in Figure 2. By training against this mixed target, SPEAR jointly learns from both sources proportionally to their signal power, rather than strictly filtering out the secondary signal. The augmented pre-training target will be used for loss computation following Eqn. (1).